find the exact value of the trigonometric function tan 5π/2 pls help! for our midterm exam

Found 2 solutions by oberobic, Alan3354:
Answer by oberobic(2304)

Using the Google calculator
.
tan((5 * pi) / 2) = 3.26635575 × 1015
.
Done

Answer by Alan3354(69443)

tan(5pi/2) = tan(pi/2) = undefined, or infinite